### v1

@@ 1 Introduction
RNA splicing is the process of removing introns from precursor mRNAs (pre-mRNAs). It is widespread in eukaryotic genomes (Keren et al., , 2010). In human, for example, each protein-coding gene contains 9.4 introns on average; $>$98% of introns start with GT on the genome (or more precisely GU on the pre-mRNA) and $>$99% end with AG (Sibley et al., , 2016). On the other hand, there are hundreds of millions of di-nucleotide GT or AG in the human genome. Only $\sim$0.1of them are real splice sites. Identifying real splice sites, which is required for gene annotation, is challenging due to the low signal-to-noise ratio.
@@
At the same time, Helixer (Holst et al., , 2023) and Tiberius (Gabriel et al., , 2024) combined Hidden Markov Models (HMM) and deep-learning models for ab initio gene prediction and achieved high accuracy. They however do not report alternative isoforms. Alignment-based methods remain the state of art for gene annotation.
While qualified spliced aligners all look for the GT..AG splice signal, they model additional flanking sequences differently. Intra-species mRNA-to-genome aligners such as BLAT (Kent, , 2002), GMAP (Wu and Watanabe, , 2005) and Splign (Kapustin et al., , 2008) often do not model extra sequences beyond GT..AG because alignment itself provides strong evidence and ambiguity shown in Fig 1 is rare. Minimap2 (Li, , 2018) prefers the GTR..YAG consensus (Irimia and Roy, , 2008). This helps to improve the alignment of noisy long RNA-seq reads. Protein-to-genome aligners tend to employ better models due to more ambiguous alignment given distant homologs. Miniprot (Li, , 2023) considers rarer GC..AG and AT..AC splice sites and optionally prioritizes on the G|GTR..YNYAG consensus common in vertebrate and insect (Iwata and Gotoh, , 2011), where “|” indicates splice boundaries. Exonerate (Slater and Birney, , 2005), Spaln (Gotoh, , 2008; Iwata and Gotoh, , 2012; Gotoh, , 2024) and the original GeneWise (Birney et al., , 2004) use PWM. GeneSeqer (Usuka and Brendel, , 2000) and GenomeThreader (Gremme et al., , 2005) apply more advanced models (Brendel and Kleffe, , 1998; Brendel et al., , 2004). Deep learning models have been applied to refining splice sites as a post-processing step (Chao et al., , 2024; Xia et al., , 2023) but have not been integrated into spliced aligners.
@@
Figure 2: Method overview. (a) Overall workflow. Training and prediction are done by minisplice. Alignment is done by minimap2 or miniprot. (b) Default model architecture and parameterization. Shaded boxes contain free parameters.
@@ 2.1 Generating training data
Due to potentially incomplete gene annotation in non-model organisms, a GT dinucleotide could be a real unannotated donor site. To alleviate this issue, we label an unannotated GT as a negative donor site only if it comes from the opposite strand of an annotated gene (Chao et al., , 2024). In the rare case when two genes on opposite strands overlap with each other, we ignore the overlapping region. To balance positive and negative sites, we randomly selsect a subset of negative sites such that the positive-to-negative ratio is 1:3. For each positive or negative donor site, we extract 100bp immediately before and after GT. The total length of sequences used for training is thus 202bp. Negative acceptor sequences are generated similarly.
@@ 2.2 Model architecture
Minisplice uses a model with two 1D-CNN layers (Fig. 2b). The architecture is common among small models for splice site prediction (Zabardast et al., , 2023). The default model uses 16 features at both CNN layers and has 7,026 free parameters in total (sum of numbers in shaded boxes). During development, we experimented alternative models, for example with different kernel sizes, more CNN or dense layers, optional dropout layers or more parameters. We chose a relatively small model in the end because it is more efficient to deploy.
@@ 2.5 Aligner integration
$\left\{\begin{array}[]{l}H_{ij}=\max\{H_{i-1,j-1}+s(i,j),E_{ij},F_{ij},\tilde{E}_{ij}-a(i)\}\\ E_{i+1,j}=\max\{H_{ij}-q,E_{ij}\}-e\\ F_{i,j+1}=\max\{H_{ij}-q,F_{ij}\}-e\\ \tilde{E}_{i+1,j}=\max\{H_{ij}-d(i)-\tilde{q},\tilde{E}_{ij}\}\\ \end{array}\right.$ (2)
@@ 3 Results
We evaluated the accuracy of trained models using Receiver Operating Characteristic (ROC) curves where we computed the true positive rate and false positive rate at different thresholds on raw model scores. To find a small model that is generalized to multiple species and is fast to deploy, we experimented models under several settings. Our final model is trained on six insect genomes from five orders and seven vertebrate genomes. We did not train a plant model because we are less familiar with the plant phylogeny.
@@
Ensembl or Gencode annotations were used for model organisms (marked by “†”); RefSeq annotations used for non-model organisms whose labels follow the naming standard developed by the Darwin Tree of Life Project: prefix “m” stands for mammals, “b” for birds, “a” for Amphibia (e.g. frogs and newts), “f” for fish, “s” for sharks, “ic” for order Coleoptera (beetles), “id” for Diptera (flies), “ih” for Hemiptera (true bugs), “il” for Lepidoptera (butterflies and moths), “iy” for Hymenoptera (bees and ants). Species marked by “*” are used for training cross-species models.
@@ 3.2 Intra-species training
Figure 3: Intra-species training. (a) Models trained on different human annotations and tested on the Gencode basic annotation. (b) Model trained from the longest human protein-coding transcripts and tested on different human annotations. (c) Donor-only training versus joint donor-acceptor training. (d) Acceptor-only training versus joint donor-acceptor training.
@@ 3.3 Cross-species training
Figure 6: Effect of window size. (a) Comparison between the default 202bp window size ($\pm$100bp around GT/AG) and 102bp. (b) Comparison between 202bp and 302bp window sizes. (c) Activation differences between positive and negative donor sites at the last max-pooling layer. With 302bp window, this layer consists of $16\times 32$ cells which are non-negative due to the ReLU function. Given a set of sequences, the activation rate of a cell is the frequency of the cell being positive (i.e. non-zero). The heat shows the rate difference between positive and negative sets. Dark red indicates more frequent activation among positive donor sites; dark blue is the opposite. (d) Activation difference for acceptor sites.
We went a step further by combining the training data across multiple vertebrate and insect species and derived model vi2, which was trained from six vertebrate and seven insect genomes marked in Table 1. Although this model is not as accurate as the species-specific model trained from individual species itself (Fig. 5a,b), it is better than applying an insect model to human (Fig. 4a). Note that vi2 is not trained on amphibian or shark genomes but it still accurately predicts splice sites in aDenEbr and sMobHyp. It is capturing common signals across large evolutionary distance while reducing overfitting to individual species. We also trained a vertebrate-only model (v2) and an insect-only model (i2). They outperformed vi2 for vertebrate and insect genomes, respectively (Fig. 5c,d), but we deemed the improvement is marginal and outweighed by the convenience of having one model across vertebrate and insect genomes.
@@
Model vi2 considers 202bp sequences around splice sites. We tried to reduce the window size to 102bp but the accuracy noticeably dropped (Fig. 6a). Increasing the window size to 302bp, on the other hand, only had minor effect (Fig. 6b). To understand what signals the model is learning, we visualized the activation rates at the last max-pooling layer which has 16 features (Fig. 6c,d). While some features such as 4 and 7 focus on signals at splice sites, other features such as 2 and 13 likely capture compositions of introns and exons. Intronic sequences around acceptors appear to provide more signals than around donors. This might be related to branch points which are located within tens of basepairs upstream to acceptors. The activation difference fades away beyond 100bp around splice sites especially within introns. It might explain why increasing window size from 202bp to 302bp has minor effect (Fig. 6b).
We further used the values at the last max-pooling layer to generate the UMAP of a random subset of training samples (Fig. 7). The UMAP separated the training samples into three large clusters which correspond to positive acceptor sites, negative sites and positive donor sites, respectively. Notably, mammals and birds contain sites absent from fish and insect. The UMAP for human, zebrafish and fruitly reveals a similar pattern. We then extracted splice sites in the UMAP that are specific to human. We found most of the intron sequences close to the splice sites tended to be short and GC-rich. We randomly selected five human genes with GC-rich donor sites, including GAPDH, ARHGEF19, ABTB2, CCAR2 and OGDH, and aligned their protein sequences to the zebrafish genome. Except CCAR2 which was aligned with low identity at 33%, other genes all reserved the GC-rich junctions, though the corresponding zebrafish junctions had lower GC.
GC-rich introns are known to exist in human (Wang and Yu, , 2011) and may help stabilize RNA secondary structure (Zhang et al., , 2011). Their specificity to mammals and birds might explain why fish and insect models do not work well for mammals (Fig. 4a) but mammalian models tend to work better for non-mammals (McCue and Burge, , 2024).
@@ 3.4 Improving spliced alignment
For the splice models, “GT-AG” only considers GT..AG, “default” considers GTR..YAG, “extended” additionally considers G|GTR..YYYAG, and “minisplice” uses the “vi2-7k” model on top of “extended”. Miniprot was invoked with “-I” which sets the maximum intron length to $\min\{\max\{3.6\sqrt{G},10^{4}\},3\times 10^{5}\}$, where $G$ is the genome size. Spaln3 was invoked with “-LS -yS -yB -yZ -yX2” and with its own splice models. %${\rm\%unannotated}=1-{\rm\#annotated}/{\rm\#predicted}$.
@@
The minisplice model also greatly reduces the junction error rate for direct RNA-seq reads (Fig. 8d), though the reduction is less pronounced when the alignment identity reaches $>$98%. For this RNA-seq run with the latest the Nanopore RNA004 kit (Zheng et al., , 2025), 74.6% of reads are mapped with identity $>$98%. The overall junction error rate is marginally reduced from 1.4% to 1.0%. The advantage of advanced splice models will be more noticeable for old data of lower quality, regions of high diversity or cross-species cDNA-to-genome alignment.
@@ 4 Discussions
If we could predict splice sites to 100% accuracy, we would be close to solving the ab initio gene finding problem. However, the conserved splice signals only come from several bases around splice sites. Given hundreds of millions of GT/AG in a mammalian genome, we would make many wrong predictions just based on the several bases. The activation pattern at an internal layer suggests our model draws power from long-range composition of exonic sequences in addition to short-range signals around splice sites. The model may tend to classify a site as donor if the sequence on the left is similar to exonic sequences in composition. With $\geq$10kb windows and orders of magnitude more parameters, recent deep learning models such as SpliceAI, Pangolin and DeltaSplice will learn composition better. They may additionally see the promoter regions of many genes and species- or even tissue-specific regulatory elements. It is not clear how much their signals come from splice sites. At this point, methods combining HMM and deep learning (Gabriel et al., , 2024; Holst et al., , 2023) may be more advantegeous for coding sequences as they explicitly model gene structures and keep protein sequences in phase.

### v2

@@ 1 Introduction
RNA splicing is the process of removing introns from precursor mRNAs (pre-mRNAs). It is widespread in eukaryotic genomes (Keren et al., , 2010). In human, for example, each protein-coding gene contains 9.4 introns on average; >>98% of introns start with GT on the genome (or more precisely GU on the pre-mRNA) and >>99% end with AG (Sibley et al., , 2016). On the other hand, there are hundreds of millions of di-nucleotide GT or AG in the human genome. Only $\sim$0.1% of them are real splice sites. Identifying real splice sites, which is required for gene annotation, is challenging due to the low signal-to-noise ratio.
@@
At the same time, Helixer (Holst et al., , 2023) and Tiberius (Gabriel et al., , 2024) combined Hidden Markov Models (HMM) and deep-learning models for ab initio gene prediction and achieved high accuracy. They however do not report alternative isoforms and are not suitable for genomes with frequent alternative splicing. Mainstream gene annotation pipelines such as Ensembl and NCBI EGAP/EGAPx still heavily rely on alignment.
While qualified spliced aligners all look for the GT..AG splice signal, they model additional flanking sequences differently. Intra-species mRNA-to-genome aligners such as BLAT (Kent, , 2002), GMAP (Wu and Watanabe, , 2005) and Splign (Kapustin et al., , 2008) often do not model extra sequences beyond GT..AG because alignment itself provides strong evidence and ambiguity shown in Fig 1 is rare. Minimap2 (Li, , 2018) prefers the GTR..YAG consensus (Irimia and Roy, , 2008). This helps to improve the alignment of noisy long RNA-seq reads. GSNAP (Wu and Nacu, , 2010) integrates MaxEnt (Yeo and Burge, , 2004) for scoring novel splice sites. Protein-to-genome aligners tend to employ better models due to more ambiguous alignment given distant homologs. Miniprot (Li, , 2023) considers rarer GC..AG and AT..AC splice sites and optionally prioritizes on the G|GTR..YNYAG consensus common in vertebrate and insect (Iwata and Gotoh, , 2011), where “|” indicates splice boundaries. Exonerate (Slater and Birney, , 2005), Spaln (Gotoh, , 2008; Iwata and Gotoh, , 2012; Gotoh, , 2024) and the original GeneWise (Birney et al., , 2004) use PWM. GeneSeqer (Usuka and Brendel, , 2000) and GenomeThreader (Gremme et al., , 2005) apply more advanced models (Brendel and Kleffe, , 1998; Brendel et al., , 2004). Deep learning models have been applied to refining splice sites as a post-processing step (Chao et al., , 2024; Xia et al., , 2023) but have not been integrated into spliced aligners.
@@
Figure 2: Method overview. (a) Overall workflow. Training and prediction are done by minisplice. Alignment is done by minimap2 or miniprot. (b) Default model architecture and parameterization, including tensor shapes. For example, $4\times 202$” indicates the input is a $4\times 202$ matrix; $16\times 4\times 5$” suggests the 1D-CNN layer has a kernel size of 5 and outputs 16 features. Shaded boxes contain free parameters.
@@ 2.1 Generating training data
Due to potentially incomplete gene annotation in non-model organisms, a GT dinucleotide could be a real unannotated donor site. To alleviate this issue, we label an unannotated GT as a negative donor site only if it comes from the opposite strand of an annotated gene (Chao et al., , 2024). In the rare case when two genes on opposite strands overlap with each other, we ignore the overlapping region. To balance positive and negative sites, we randomly selsect a subset of negative sites such that the positive-to-negative ratio is 1:3. For each positive or negative donor site, we extract 100bp immediately before and after GT. The total length of sequences used for training is thus 202bp. Negative acceptor sequences are generated similarly. We also experimented 102bp and 302bp window sizes.
@@ 2.2 Model architecture
Minisplice uses a model with two 1D-CNN layers (Fig. 2b). The architecture is common among small models for splice site prediction (Zabardast et al., , 2023). The default model uses 16 features at both CNN layers and has 7,026 free parameters in total (sum of numbers in shaded boxes). During development, we experimented alternative models with different kernel sizes, more CNN or dense layers, optional dropout layers or more parameters. We chose a relatively small model in the end as it is more efficient to deploy. Because minimap2 and miniprot score each donor or acceptor site independently, we also model splice sites independently.
@@ 2.5 Aligner integration
[…]
@@ 3 Results
We evaluated the accuracy of trained models using Receiver Operating Characteristic (ROC) curves where we computed the true positive rate and false positive rate at different thresholds on raw model scores. In the ROC plot, we focused on the region with sensitivity above 50% and false positive rate below 10% because we intend to improve spliced alignment in this region. For each curve, we calculated rAUC, which is the area the under the ROC curve restricted to and scaled by this region. Due to scaling, ${\rm rAUC}\in[0,1]$.
To find a small model that is generalized to multiple species and is fast to deploy, we experimented models under several settings. Our final model is trained on six insect genomes from five orders and seven vertebrate genomes. We did not train a plant model because we are less familiar with the plant phylogeny.
@@
bTaeGut Taeniopygia guttata GCF_048771995.1
rEmyOrb Emys orbicularis GCF_028017835.1
@@
Ensembl or Gencode annotations were used for model organisms (marked by “†”); RefSeq annotations used for non-model organisms whose labels follow the naming standard developed by the Darwin Tree of Life Project: prefix “m” stands for mammals, “b” for birds, “r” for reptiles, “a” for amphibians, “f” for fish, “s” for sharks, “ic” for order Coleoptera (beetles), “id” for Diptera (flies), “ih” for Hemiptera (true bugs), “il” for Lepidoptera (butterflies and moths), “iy” for Hymenoptera (bees and ants). Species marked by “*” are used for training cross-species models.
@@ 3.2 Intra-species training
Figure 3: Intra-species training. (a) Models trained on different human annotations and tested on the Gencode basic annotation. Numbers in parentheses denote rAUC, which is the Area Under the Curve restricted to the plotted region. (b) Model trained from the longest human protein-coding transcripts and tested on different human annotations. (c) Donor-only training versus joint donor-acceptor training. (d) Acceptor-only training versus joint donor-acceptor training.
@@ 3.3 Cross-species training
Figure 6: Effect of window size. (a) Comparison between the default 202bp window size ($\pm$100bp around GT/AG) and 102bp. (b) Comparison between 202bp and 302bp window sizes. (c) Activation differences between positive donor sites and negative GT sites at the last max-pooling layer. Positive coordinates correspond to exonic bases. With 302bp window, this layer consists of $16\times 32$ cells which are non-negative due to the ReLU function. Given a set of sequences, the activation rate of a cell is the frequency of the cell being positive (i.e. non-zero). The heat shows the rate difference between positive and negative sets. Dark red indicates more frequent activation among positive donor sites; dark blue is the opposite. (d) Activation difference for acceptor sites.
We went a step further by combining the training data across multiple vertebrate and insect species and derived model vi2, which was trained from six vertebrate and seven insect genomes marked in Table 1. Although this model is not as accurate as the species-specific model trained from individual species itself (Fig. 5a,b), it is better than applying an insect model to human (Fig. 4a). Note that vi2 is not trained on amphibian or shark genomes but it still accurately predicts splice sites in aDenEbr and sMobHyp. It is capturing common signals across large evolutionary distance while reducing overfitting to individual species. We also trained a vertebrate-only model (v2) and an insect-only model (i2). They outperformed vi2 for vertebrate and insect genomes, respectively (Fig. 5c,d), but we deemed the improvement is small and outweighed by the convenience of having one model across vertebrate and insect genomes.
@@
Figure 8: GC content of homologous introns. Vertebrate protein sequences from Swiss-Prot are aligned to each genome with miniprot. Given two genomes, if a protein sequence is aligned with introns at the same position on the protein, the two aligned introns are considered homologous to each other. (a) Comparing percent GC content between homologous human and zebrafish introns. (b) Human versus hourglass tree frog (aDenEbr). (c) Human versus zebra finch (bTaeGut). (d) Zebra finch versus European pond turtle (rEmyOrb).
Model vi2 considers 202bp sequences around splice sites. We tried to reduce the window size to 102bp but the accuracy dropped (Fig. 6a). Increasing the window size to 302bp, on the other hand, only had a small effect (Fig. 6b). To understand what signals the model is learning, we visualized the activation rates at the last max-pooling layer which has 16 features (Fig. 6c,d). While some features such as 4 and 7 focus on signals at splice sites, other features such as 2 and 13 likely capture compositions of introns and exons. Intronic sequences around acceptors appear to provide more signals than around donors. This might be related to branch points which are located within tens of basepairs upstream to acceptors. The activation difference fades away beyond 100bp around splice sites especially within introns. It is possibly why increasing window size from 202bp to 302bp has minor effect (Fig. 6b).
We further used the values at the last max-pooling layer to generate the UMAP of a random subset of training samples (Fig. 7). The UMAP separated the training samples into three large clusters which correspond to positive acceptor sites, negative sites and positive donor sites, respectively. Notably, some areas in the UMAP (the two red circles) only contain splice sites from mammals and birds but rarely from fish and insect. We extracted splice sites in these areas and found most of them came from GC-rich introns. The presence of these areas might explain why fish and insect models do not work well for mammals (Fig. 4a) but mammalian models tend to work better for non-mammals (McCue and Burge, , 2024).
To more directly investigate these GC-rich introns, we extracted homologous intron pairs and compared their GC content (Fig. 8). We saw introns reaching 60% GC in human but their homologs in zebrafish are <<40% in GC (Fig. 8a). Consistent with our earlier observation in UMAP, few zebrafish introns are GC-rich. Although the GC content is overall higher in amphibians, not many introns can reach 60% GC (Fig. 8b). Birds have high-GC introns like human (Fig. 8c). This is unrelated to warm- versus cold-blooded organisms per isochore theory (Bernardi et al., , 1985) because reptiles also have GC-rich introns (Fig. 8d).
GC-rich introns are known to exist in human (Wang and Yu, , 2011) and may help stabilize RNA secondary structure (Zhang et al., , 2011). They are more likely to be retained than GC-poor introns (Sibley et al., , 2016), and they tend to be located in the nuclear center (Choquet et al., , 2025) which may be partly correlated with the higher CpG dentity in the center (Xie et al., , 2017; Tan et al., , 2018).
@@ 3.4 Improving spliced alignment
For the splice models, “GT-AG” only considers GT..AG, “default” considers GTR..YAG, “extended” additionally considers G|GTR..YYYAG, and “minisplice” uses the “vi2-7k” model on top of “extended”. Miniprot was invoked with “-I” which sets the maximum intron length to $\min\{\max\{3.6\sqrt{G},10^{4}\},3\times 10^{5}\}$, where $G$ is the genome size. Spaln3 was invoked with “-LS -yS -yB -yZ -yX2” and with its own splice models. ${\rm\%unannotated}=1-{\rm\#annotated}/{\rm\#predicted}$.
@@
We were evaluating junction accuracy. The exon accuracy was 9–11% lower for the zebrafish-to-human alignment. We manually inspected some exon-specific errors and classified them into three cases. First, an alignment did not reach the start of the protein sequence. The first junction in the alignment could still be correct but the beginning of the first exon was often an error. Second, 10.7% of longest zebrafish proteins did not start with methionine “M”, the start amino acid. Third, zebrafish proteins starting with “M” may occasionally be mutated in human. This is an interesting but rare case. Overall, we intend to improve junction accuracy in this work but not the three cases above. Junction accuracy serves our goal better.
The minisplice model also greatly reduces the junction error rate for direct RNA-seq reads (Fig. 9d), though the reduction is less pronounced when the alignment identity reaches >>98%. For this RNA-seq run with the latest the Nanopore RNA004 kit (Zheng et al., , 2025), 74.6% of reads are mapped with identity >>98%. The overall junction error rate is marginally reduced from 1.4% to 1.0%. The advantage of advanced splice models will be more noticeable for old data of lower quality, regions of high diversity or cross-species cDNA-to-genome alignment.
@@
As to performance, minisplice precomputed splice scores for the human genome in 1.4 hours over 16 CPU threads. This step only needs to be done once per genome. Minimap2 and miniprot took $\sim$13 seconds upfront to load precomputed splice scores in the gzip format. The following alignment speed was barely affected as looking up splice scores is much faster than alignment.
@@ 4 Discussions
If we could predict splice sites to 100% accuracy, we would be close to solving the ab initio gene finding problem. However, the conserved splice signals only come from several bases around splice sites. Given hundreds of millions of GT/AG in a mammalian genome, we would make many wrong predictions just based on the several bases. The activation pattern at an internal layer suggests our model draws power from long-range composition of exonic sequences in addition to short-range signals around splice sites. The model may tend to classify a site as donor if the sequence on the left is similar to exonic sequences in composition. Nonetheless, even at 1% false positive rate, our model would predict 3.75 million ($3\times 10^{9}\times 2/16\times 1\%$) false donor sites on average, an order of magnitude more than real donor sites. We must consider additional information for more accurate prediction. With $\geq$10kb windows and orders of magnitude more parameters, recent deep learning models such as SpliceAI, Pangolin and DeltaSplice will learn composition better. They may additionally see the promoter regions of many genes and species- or even tissue-specific regulatory elements. It is not clear how much their signals come from splice sites. At this point, methods combining HMM and deep learning (Gabriel et al., , 2024; Holst et al., , 2023) may be more advantageous for coding sequences as they explicitly model gene structures and keep protein sequences in phase.